Use iteration to guess the formula:

di-3dk-i +1 withk2 1 do 3
Answer #1
dk = 3(dk-1) + 1
= 3(3(dk-2) + 1) + 1
= 3(3(3(dk-3) + 1) + 1) + 1
= 3^3 dk-3 + 3^2+3^1+3^0
= 3^k d0 + 3^k-1+...+3^1+3^0
= 3^k*3 + (3^k - 1)/2
= (3^k*6 + 3^k - 1)/2
= (7 3^k - 1)/2
so, dk is (7*3^k - 1)/2